OpenStudy (anonymous):

What is the solution to the rational equation x/2x-1 - 1/4 = 2/2x-1 ?

OpenStudy (dtan5457):

Possibly my worst topic to date. Someone teach us, lol.

OpenStudy (mathmate):

Rewrite it as \(\large\frac{x}{2x-1}-\frac{2}{2x-1}=\frac{1}{4}\)

OpenStudy (mathmate):

Next step: subtract and then cross multiply.

OpenStudy (anonymous):

so 2x^2-1 and 4x-1 ?

OpenStudy (dtan5457):

@freckles

OpenStudy (dtan5457):

can you finish this?

OpenStudy (freckles):

it looks like mathmate is here to your rescues i will be a cheerleader for now and root for mathmate

OpenStudy (mathmate):

If you noticed the common denominator on the left hand side that would make an easy subtraction: \(\large\frac{x-2}{2x-1}=\frac{1}{4}\) All you have left is cross multiply and solve for x.

OpenStudy (mathmate):

@freckles Everybody can join in! Get wet!

OpenStudy (freckles):

Check your solutions to see if they are within the domain of the equation

OpenStudy (anonymous):

would it be 7/2?

OpenStudy (dtan5457):

ohh. I get it now. and yes it's 7/2 ^^

OpenStudy (freckles):

you could have also chose to multiply your original equation by 2x-1 to get rid of the denominator (while keeping in mind that x cannot be 1/2) \[x-\frac{1}{4}(2x-1)=2 \] if you don't like that 4 in the bottom multiply both sides by 4 you have 4x-(2x-1)=8 now solve for x 4x-2x+1=8 2x+1=8 2x=7 x=7/2

OpenStudy (anonymous):

What is the simplified form of x^2 - 5x + 6/15xy^2 all over 2x^2 - 7x + 3 /5x^2y? could you help with this one also?

OpenStudy (freckles):

\[\frac{\frac{x^2-5x+6}{15xy^2}}{\frac{2x^2-7x+3}{5x^2y}} \text{ or read \it this way } \\ \frac{x^2-5x+6}{15xy^2} \cdot \frac{5x^2y}{2x^2-7x+3}\] do some factoring and see if things cancel